OpenStudy (anonymous):

What is the equation of the line that passes through the points (0, -4) and (2, 8)? y= 6x- 4 y= 6x+ 4 y= 4x+ 6 y= -4x+ 6

OpenStudy (anonymous):

first find the slope then use point slope form \[\Large m=\frac{8-(-4)}{2-0}=\frac{12}{2}=6\] now it is upto you use any point from the above two and use this formula (point slope form) \[\Large y-y_{0}=m(x-x_{0})\] m is already found and (x0,y0) is any point you use to use

OpenStudy (anonymous):

@miadeex can you do this now?

OpenStudy (anonymous):

any point you choose to use !

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

ok it should be a

OpenStudy (anonymous):

remember it is -4 when using the above equation !

OpenStudy (waleed_imtiaz):

Use slope formula to calculate the slope and then apply point-slope form..... slope= y2-y1/x2-x1 Which gives.... 6 now, put any of the point... in the point slope form.... U will get...... y-8=6(x-2) then ...y= 6x-4 Option A